# HEAL: An Empirical Study on in mbodied gents Driven by arge Language Models

### v1

@@ Abstract
Trishna ChakrabortyāUdita Ghosh āXiaopan ZhangāFahim Faisal Niloy Yue Dong āJiachen Li āAmit K. Roy-Chowdhury āChengyu Song University of California, Riverside {tchak006,ughos002,xzhan006,fnilo001,jiachen.li,yue.dong}@ucr.edu csong@cs.ucr.edu āamitrc@ece.ucr.edu
@@ 1 Introduction
Because existing datasets either do not explicitly aim to elicit hallucination in long-horizon embodied tasks, or do not target embodied agent tasks, we first construct a new probing set that is more likely to cause hallucination. Constructing such a probing set faces two main challenges: (i) how to induce hallucinations and (ii) how to detect unwanted behaviors caused by hallucinations. We solve these challenges by systematically modifying base prompts from the existing LLM-based embodied agent benchmarkĀ Li etĀ al. (2024b) to introduce scene-task inconsistencies. Specifically, we modify the two core components of each base promptāthe task description and the scene informationāto create mismatches between user instructions and the observed environment (see FigureĀ 2).
@@ 2 Related Work
Hallucinations in LLMs. Hallucinations in LLMs have been widely studied in text-onlyĀ Dhuliawala etĀ al. (2023); Agrawal etĀ al. (2024). Recent works extend these to multimodal LLMs Ā Bai etĀ al. (2024); Xu etĀ al. (2025); Guan etĀ al. (2024). Causes include data quality issues, weak vision encoders, and language model priors overriding visual evidence. Mitigation approaches include cross-checking Ā Yu etĀ al. (2024b), instruction-tuningĀ Liu etĀ al. (2023), self-correctionĀ Peng etĀ al. (2023), and others.
@@ 3.2 Hallucination Instances
Env | Models | Base Prompt | TaskDescMod | SceneMod | |||||||||||
DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||||
V I R T U A L H O M E | Llama-3-8B-Instruct | 3.7 | 1.2 | 25.6 | 4.8 | 60.51 | 20.4 | 3.7 | 36.60 | 17.0 | 2.6 | 27.36 | 38.5 | 8.7 | 66.45 |
DS-R1-Distil-LLaMA-8B | 19.7 | 4.7 | 48.6 | 7.5 | 56.85 | 37.6 | 5.7 | 38.49 | 40.2 | 5.9 | 30.78 | 57.1 | 9.3 | 63.36 | |
Gemma-2-9b-it | 4.4 | 5.4 | 27.1 | 9.9 | 81.02 | 31.7 | 8.7 | 46.56 | 16.2 | 6.1 | 28.34 | 52.6 | 15.0 | 66.78 | |
Qwen-14B-Instruct | 3.6 | 4.2 | 26.4 | 4.4 | 72.69 | 36.6 | 6.0 | 47.59 | 13.1 | 4.8 | 17.10 | 50.6 | 8.0 | 70.68 | |
DS-R1-Distil-Qwen-14B | 5.6 | 2.9 | 28.2 | 3.7 | 50.76 | 33.4 | 5.0 | 39.86 | 14.6 | 3.9 | 26.22 | 56.5 | 7.3 | 69.87 | |
Llama-4-Scout-17B-16E-Instruct | 1.6 | 0.7 | 6.3 | 3.3 | 18.88 | 21.1 | 9.2 | 36.25 | 14.3 | 1.3 | 26.38 | 48.2 | 8.9 | 67.26 | |
Llama-3.3-70B-Instruct | 1.5 | 0.2 | 12.8 | 0.2 | 21.02 | 22.8 | 1.9 | 46.74 | 8.0 | 1.0 | 21.82 | 37.6 | 0.8 | 60.10 | |
Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 8.27 | 3.0 | 0.0 | 7.21 | 6.0 | 0.0 | 20.08 | |
Gemini 2.5 Flash | 0.5 | 0.0 | 0.6 | 0.0 | 0.2 | 22.4 | 0.0 | 50.86 | 5.7 | 0.0 | 7.82 | 36.0 | 0.0 | 61.89 | |
Gemini 2.5 Flash (w/o thinking) | 1.3 | 0.0 | 2.1 | 1.0 | 5.18 | 14.4 | 0.8 | 39.00 | 12.7 | 0.0 | 20.03 | 18.4 | 0.0 | 47.07 | |
Mistral-Large-Instruct-2411 | 1.6 | 1.0 | 14.0 | 1.4 | 8.73 | 21.1 | 1.5 | 54.64 | 15.2 | 1.6 | 17.10 | 47.1 | 2.1 | 67.75 | |
Claude 3.5 Sonnet | 2.1 | 0.1 | 2.2 | 0.3 | 0.2 | 16.6 | 0.2 | 45.53 | 31.0 | 0.0 | 36.64 | 14.4 | 1.0 | 40.07 | |
GPT-4o | 1.0 | 1.2 | 2.9 | 2.0 | 3.45 | 7.6 | 1.8 | 31.10 | 22.6 | 1.3 | 27.04 | 36.8 | 1.4 | 60.91 | |
B E H A V I O R | Llama-3-8B-Instruct | 1.93 | 2.15 | 8.2 | 10.8 | 6.40 | 5.0 | 6.7 | 14.34 | 2.2 | 7.4 | 0.70 | 17.8 | 7.7 | 18.27 |
DS-R1-Distil-LLaMA-8B | 1.26 | 8.16 | 13.9 | 27.2 | 25.25 | 6.8 | 13.3 | 28.84 | 20.4 | 10.1 | 21.39 | 60 | 19.6 | 25.65 | |
Gemma-2-9b-it | 0.87 | 1.8 | 11.6 | 5.7 | 9.43 | 10.8 | 4.8 | 22.80 | 11.4 | 2.8 | 5.74 | 73.1 | 8.2 | 19.74 | |
Qwen-14B-Instruct | 0.41 | 3.33 | 15.8 | 5.7 | 13.13 | 6.6 | 4.6 | 18.48 | 1.7 | 4.1 | 1.57 | 63.1 | 5.7 | 20.76 | |
DS-R1-Distil-Qwen-14B | 0.32 | 1.24 | 8.4 | 5.2 | 13.13 | 4.5 | 6.0 | 19.17 | 2.8 | 7.6 | 4.52 | 46.4 | 4.7 | 17.99 | |
Llama-4-Scout-17B-16E-Instruct | 0.0 | 1.7 | 0.3 | 2.3 | 0.34 | 2.6 | 1.7 | 11.40 | 3.9 | 1.0 | 3.48 | 32.0 | 1.7 | 15.22 | |
Llama-3.3-70B-Instruct | 0.0 | 2.1 | 1.9 | 2.8 | 2.69 | 3.2 | 1.7 | 15.89 | 0.0 | 1.8 | 0.0 | 45.0 | 1.3 | 12.82 | |
Gemini 2.0 Flash | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 1.1 | 0.2 | 4.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.6 | 4.84 | 0.0 | 0.0 | 0.0 | 12.6 | 1.1 | 5.54 | |
Gemini 2.5 Flash (w/o thinking) | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.4 | 1.4 | 6.22 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | |
Mistral-Large-Instruct-2411 | 0.0 | 0.0 | 0.2 | 1.6 | 0.34 | 2.8 | 0.4 | 12.95 | 0.0 | 0.5 | 0.0 | 38.6 | 0.0 | 12.55 | |
Claude 3.5 Sonnet | 0.0 | 2.0 | 0.0 | 3.0 | 0.0 | 1.0 | 1.2 | 3.28 | 0.0 | 2.5 | 0.0 | 0.0 | 2.8 | 0.0 | |
GPT-4o | 0.0 | 0.0 | 1.9 | 0.5 | 2.36 | 2.6 | 0.0 | 10.71 | 0.0 | 0.0 | 0.0 | 31.9 | 0.0 | 9.87 |
@@ 4.1 Experimental Setup
Datasets. We evaluate long-horizon tasks that require multiple sequential steps across two simulation environments: VirtualHomeĀ Puig etĀ al. (2018) and BEHAVIORĀ Li etĀ al. (2023a). TableĀ 11 in the Appendix shows the prompt distribution in our probing set, which contains 2,574 samples designed to demonstrate task-scene inconsistencies.
Models. We evaluate 12 open and closed LLMs, including models from LLaMAĀ Grattafiori etĀ al. (2024), QwenĀ Yang etĀ al. (2024a), GemmaĀ Team etĀ al. (2024), GeminiĀ Team etĀ al. (2023), ClaudeĀ Anthropic (2024), MistralĀ AI (2024), GPT-4oĀ Achiam etĀ al. (2023), as well as DeepSeek-R1Ā Guo etĀ al. (2025) distilled versions of LLaMA and Qwen. To assess the role of vision, we evaluate VLMs with the same language backbones, including Qwen-VLĀ Wang etĀ al. (2024a), Gemma-3Ā Team etĀ al. (2025), and LLaMA-Vision. Smaller models are more suitable for running locally in robots, while large models represent the state-of-the-art. See TableĀ 12 for model cards.
@@ 4.2 Results
Models | Mitigation Method | TaskDescMod | SceneMod | ||||||||||
DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||
Llama-3-8B-Instruct | KAF | 9.0 | 1.4 | 45.99 | 10.9 | 2.8 | 30.93 | 7.9 | 1.6 | 24.76 | 19.4 | 3.3 | 59.12 |
SCW | 2.3 | 1.5 | 0.0 | 7.9 | 2.6 | 5.15 | 4.8 | 1.8 | 7.33 | 17.2 | 5.4 | 25.41 | |
Gemma-2-9b-it | KAF | 10.3 | 5.3 | 65.48 | 17.6 | 8.5 | 42.96 | 9.9 | 4.1 | 25.08 | 38.8 | 9.9 | 60.59 |
SCW | 4.4 | 6.5 | 5.28 | 11.5 | 8.6 | 19.59 | 6.7 | 5.3 | 12.05 | 38.2 | 10.9 | 34.69 | |
Qwen-14B-Instruct | KAF | 12.1 | 2.8 | 47.72 | 21.4 | 5.0 | 41.24 | 5.3 | 2.2 | 10.10 | 18.9 | 4.9 | 55.37 |
SCW | 2.0 | 2.6 | 0.30 | 18.7 | 6 | 17.35 | 5.0 | 3.3 | 3.75 | 20.8 | 4.2 | 42.35 |
@@
State Hallucination. State hallucinations stay relatively low, consistent with the fact that we do not explicitly introduce state inconsistencies. A moderate increase in state hallucination compared to the base prompts indicates secondary effects: object hallucinations may also trigger incorrect state predictions. See AppendixĀ F for examples.
@@ Appendix E Toward Reliable Planning
Appendix F State Hallucinations
@@
Appendix G Hallucination Probing Prompts Design and Distribution
G.1 Base Prompt Format and Structure

### v2

@@ Abstract
Trishna ChakrabortyāāUdita Ghosh āāXiaopan ZhangāāFahim Faisal Niloy Yue Dong āJiachen Li āAmit K. Roy-Chowdhury āChengyu Song University of California, Riverside {tchak006, ughos002, xzhan006, fnilo001, yue.dong, jiachen.li}@ucr.edu āamitrc@ece.ucr.edu ācsong@cs.ucr.edu
@@ 1 Introduction
Because existing datasets either do not explicitly aim to elicit hallucination in long-horizon embodied tasks, or do not target embodied agent tasks, we first construct a new probing set111The HEAL probing set is publicly available at https://huggingface.co/datasets/Trishna13/HEAL that is more likely to cause hallucination. Constructing such a probing set faces two main challenges: (i) how to induce hallucinations and (ii) how to detect unwanted behaviors caused by hallucinations. We solve these challenges by systematically modifying base prompts from the existing LLM-based embodied agent benchmarkĀ Li etĀ al. (2024b) to introduce scene-task inconsistencies. Specifically, we modify the two core components of each base promptāthe task description and the scene informationāto create mismatches between user instructions and the observed environment (see FigureĀ 2).
@@ 2 Related Work
Hallucinations in LLMs. Hallucinations in LLMs have been widely studied in text-only settingsĀ Dhuliawala etĀ al. (2023); Agrawal etĀ al. (2024). Recent work extends these to multimodal LLMs, manifested as category misidentification, attribute errors, or relation misrepresentation Ā Bai etĀ al. (2024); Xu etĀ al. (2025); Guan etĀ al. (2024). Causes include data quality issues, weak vision encoders, and language model priors overriding visual evidence. Mitigation approaches include cross-checking Ā Yu etĀ al. (2024b), instruction-tuningĀ Liu etĀ al. (2023), self-correctionĀ Peng etĀ al. (2023), and others.
@@ 3.2 Hallucination Instances
| Env | Models | Base Prompt | TaskDescMod | SceneMod | ||||||||||||||||||||||
| DistInj | TaskObjRem | SynonymSub | SceneTaskCon | |||||||||||||||||||||||
Llama-3-8B-Instruct | 3.7 | 1.2 | 25.6 | 4.8 | 60.51 | 20.4 | 3.7 | 36.60 | 17.0 | 2.6 | 27.36 | 38.5 | 8.7 | 66.45 | |||||||||||
| DS-R1-Distil-LLaMA-8B | 19.7 | 4.7 | 48.6 | 7.5 | 56.85 | 37.6 | 5.7 | 38.49 | 40.2 | 5.9 | 30.78 | 57.1 | 9.3 | 63.36 | ||||||||||||
| Gemma-2-9b-it | 4.4 | 5.4 | 27.1 | 9.9 | 81.02 | 31.7 | 8.7 | 46.56 | 16.2 | 6.1 | 28.34 | 52.6 | 15.0 | 66.78 | ||||||||||||
| Qwen-14B-Instruct | 3.6 | 4.2 | 26.4 | 4.4 | 72.69 | 36.6 | 6.0 | 47.59 | 13.1 | 4.8 | 17.10 | 50.6 | 8.0 | 70.68 | ||||||||||||
| DS-R1-Distil-Qwen-14B | 5.6 | 2.9 | 28.2 | 3.7 | 50.76 | 33.4 | 5.0 | 39.86 | 14.6 | 3.9 | 26.22 | 56.5 | 7.3 | 69.87 | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | 1.6 | 0.7 | 6.3 | 3.3 | 18.88 | 21.1 | 9.2 | 36.25 | 14.3 | 1.3 | 26.38 | 48.2 | 8.9 | 67.26 | ||||||||||||
| Llama-3.3-70B-Instruct | 1.5 | 0.2 | 12.8 | 0.2 | 21.02 | 22.8 | 1.9 | 46.74 | 8.0 | 1.0 | 21.82 | 37.6 | 0.8 | 60.10 | ||||||||||||
| Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 8.27 | 3.0 | 0.0 | 7.21 | 6.0 | 0.0 | 20.08 | ||||||||||||
| Gemini 2.5 Flash | 0.5 | 0.0 | 0.6 | 0.0 | 0.2 | 22.4 | 0.0 | 50.86 | 5.7 | 0.0 | 7.82 | 36.0 | 0.0 | 61.89 | ||||||||||||
| Gemini 2.5 Flash (w/o thinking) | 1.3 | 0.0 | 2.1 | 1.0 | 5.18 | 14.4 | 0.8 | 39.00 | 12.7 | 0.0 | 20.03 | 18.4 | 0.0 | 47.07 | ||||||||||||
| Mistral-Large-Instruct-2411 | 1.6 | 1.0 | 14.0 | 1.4 | 8.73 | 21.1 | 1.5 | 54.64 | 15.2 | 1.6 | 17.10 | 47.1 | 2.1 | 67.75 | ||||||||||||
| Claude 3.5 Sonnet | 2.1 | 0.1 | 2.2 | 0.3 | 0.2 | 16.6 | 0.2 | 45.53 | 31.0 | 0.0 | 36.64 | 14.4 | 1.0 | 40.07 | ||||||||||||
| GPT-4o | 1.0 | 1.2 | 2.9 | 2.0 | 3.45 | 7.6 | 1.8 | 31.10 | 22.6 | 1.3 | 27.04 | 36.8 | 1.4 | 60.91 | ||||||||||||
Llama-3-8B-Instruct | 1.93 | 2.15 | 8.2 | 10.8 | 6.40 | 5.0 | 6.7 | 14.34 | 2.2 | 7.4 | 0.70 | 17.8 | 7.7 | 18.27 | |||||||||||
| DS-R1-Distil-LLaMA-8B | 1.26 | 8.16 | 13.9 | 27.2 | 25.25 | 6.8 | 13.3 | 28.84 | 20.4 | 10.1 | 21.39 | 60 | 19.6 | 25.65 | ||||||||||||
| Gemma-2-9b-it | 0.87 | 1.8 | 11.6 | 5.7 | 9.43 | 10.8 | 4.8 | 22.80 | 11.4 | 2.8 | 5.74 | 73.1 | 8.2 | 19.74 | ||||||||||||
| Qwen-14B-Instruct | 0.41 | 3.33 | 15.8 | 5.7 | 13.13 | 6.6 | 4.6 | 18.48 | 1.7 | 4.1 | 1.57 | 63.1 | 5.7 | 20.76 | ||||||||||||
| DS-R1-Distil-Qwen-14B | 0.32 | 1.24 | 8.4 | 5.2 | 13.13 | 4.5 | 6.0 | 19.17 | 2.8 | 7.6 | 4.52 | 46.4 | 4.7 | 17.99 | ||||||||||||
| Llama-4-Scout-17B-16E-Instruct | 0.0 | 1.7 | 0.3 | 2.3 | 0.34 | 2.6 | 1.7 | 11.40 | 3.9 | 1.0 | 3.48 | 32.0 | 1.7 | 15.22 | ||||||||||||
| Llama-3.3-70B-Instruct | 0.0 | 2.1 | 1.9 | 2.8 | 2.69 | 3.2 | 1.7 | 15.89 | 0.0 | 1.8 | 0.0 | 45.0 | 1.3 | 12.82 | ||||||||||||
| Gemini 2.0 Flash | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 1.1 | 0.2 | 4.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||||||||
| Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.6 | 4.84 | 0.0 | 0.0 | 0.0 | 12.6 | 1.1 | 5.54 | ||||||||||||
| Gemini 2.5 Flash (w/o thinking) | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.4 | 1.4 | 6.22 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | ||||||||||||
| Mistral-Large-Instruct-2411 | 0.0 | 0.0 | 0.2 | 1.6 | 0.34 | 2.8 | 0.4 | 12.95 | 0.0 | 0.5 | 0.0 | 38.6 | 0.0 | 12.55 | ||||||||||||
| Claude 3.5 Sonnet | 0.0 | 2.0 | 0.0 | 3.0 | 0.0 | 1.0 | 1.2 | 3.28 | 0.0 | 2.5 | 0.0 | 0.0 | 2.8 | 0.0 | ||||||||||||
| GPT-4o | 0.0 | 0.0 | 1.9 | 0.5 | 2.36 | 2.6 | 0.0 | 10.71 | 0.0 | 0.0 | 0.0 | 31.9 | 0.0 | 9.87 | ||||||||||||
@@ 4.1 Experimental Setup
Datasets. We evaluate long-horizon tasks that require multiple sequential steps across two simulation environments: VirtualHomeĀ Puig etĀ al. (2018) and BEHAVIORĀ Li etĀ al. (2023a). TableĀ 13 in the Appendix shows the prompt distribution in our probing set, which contains 2,574 samples designed to demonstrate task-scene inconsistencies.
Models. We evaluate 12 open and closed LLMs, including models from LLaMAĀ Grattafiori etĀ al. (2024), QwenĀ Yang etĀ al. (2024a), GemmaĀ Team etĀ al. (2024), GeminiĀ Team etĀ al. (2023), ClaudeĀ Anthropic (2024), MistralĀ AI (2024), GPT-4oĀ Achiam etĀ al. (2023), as well as DeepSeek-R1Ā Guo etĀ al. (2025) distilled versions of LLaMA and Qwen. To assess the role of vision, we evaluate VLMs with the same language backbones, including Qwen-VLĀ Wang etĀ al. (2024a), Gemma-3Ā Team etĀ al. (2025), and LLaMA-Vision. Smaller models are more suitable for running locally in robots, while large models represent the state-of-the-art. All experiments are performed with three independent runs per setting, and results reflect the mean performance. See TableĀ 14 for model cards.
@@ 4.2 Results
| Models |
TaskDescMod | SceneMod | ||||||||||||
| DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||||
| Llama-3-8B-Instruct | KAF | 9.0 | 1.4 | 45.99 | 10.9 | 2.8 | 30.93 | 7.9 | 1.6 | 24.76 | 19.4 | 3.3 | 59.12 | ||
| SCW | 2.3 | 1.5 | 0.0 | 7.9 | 2.6 | 5.15 | 4.8 | 1.8 | 7.33 | 17.2 | 5.4 | 25.41 | |||
| Gemma-2-9b-it | KAF | 10.3 | 5.3 | 65.48 | 17.6 | 8.5 | 42.96 | 9.9 | 4.1 | 25.08 | 38.8 | 9.9 | 60.59 | ||
| SCW | 4.4 | 6.5 | 5.28 | 11.5 | 8.6 | 19.59 | 6.7 | 5.3 | 12.05 | 38.2 | 10.9 | 34.69 | |||
| Qwen-14B-Instruct | KAF | 12.1 | 2.8 | 47.72 | 21.4 | 5.0 | 41.24 | 5.3 | 2.2 | 10.10 | 18.9 | 4.9 | 55.37 | ||
| SCW | 2.0 | 2.6 | 0.30 | 18.7 | 6 | 17.35 | 5.0 | 3.3 | 3.75 | 20.8 | 4.2 | 42.35 | |||
@@
State Hallucination. State hallucinations stay relatively low, consistent with the fact that we do not explicitly introduce state inconsistencies. A moderate increase in state hallucination compared to the base prompts indicates secondary effects: object hallucinations may also trigger incorrect state predictions. See subsectionĀ F.1 for examples.
@@ 7 Limitations
Acknowledgments
This research was supported and funded by the NSF grants CMMI-2326309, III-1901379, CNS-2312395, 2431569, UC MRPI and SoCalHub. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein. Finally, we are grateful to the anonymous reviewers for their constructive feedback.
@@ Appendix E Toward Reliable Planning
Appendix F Other Forms of Hallucinations
F.1 State Hallucinations
@@
F.2 Relation Hallucinations
Following our definition in subsectionĀ 3.1, relation hallucination occurs when the predicted relations fall outside the allowable set. As shown in TableĀ 11, smaller models exhibit higher rates of relation hallucination, consistent with trends observed for object and state hallucinations in the main paper. We highlight that the relation hallucinations observed here primarily arise as a secondary effect of the object-level inconsistencies introduced in our study. Systematically designing scenarios to explicitly induce attribute- or spatial-level hallucinations remains a valuable but distinct direction, which we leave for future work.
| Models | Base Prompt | DistInj | TaskObjRem | SynonymSub | SceneTaskCon |
| Llama-3-8B-Instruct | 17.0 | 22.9 | 18.5 | 17.3 | 19.0 |
| DS-R1-Distil-LLaMA-8B | 19.6 | 30.2 | 19.8 | 23.4 | 23.4 |
| Gemma-2-9b-it | 3.8 | 4.8 | 8.0 | 6.5 | 3.9 |
| Qwen-14B-Instruct | 13.1 | 14.5 | 19.9 | 24.7 | 13.9 |
| DS-R1-Distil-Qwen-14B | 10.9 | 15.4 | 13.3 | 14.9 | 11.1 |
| Llama-4-Scout-17B-16E-Instruct | 6.6 | 10.6 | 6.8 | 7.1 | 7.7 |
| Llama-3.3-70B-Instruct | 0.0 | 0.9 | 0.0 | 0.3 | 0.4 |
| Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Mistral-Large-Instruct-2411 | 0.3 | 0.8 | 0.4 | 1.6 | 0.5 |
| Claude 3.5 Sonnet | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| GPT-4o | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
Appendix G Hallucinations on Core vs. Peripheral Object Removal
We qualitatively observe that hallucinations are more frequent when core objects are removed. TableĀ 12 presents case-by-case examples where we remove one object at a time and report the corresponding hallucination rates. The results show that removing essential objects, such as the pot while cooking, the dishwasher while washing dishes, or the washing machine while doing laundry, leads to high hallucination rates, while removing peripheral objects has little to no effect.
| Task Name | Corresponding Scene Objects Removed (one at a time) | Model Name | Hallucination Rate |
| Cook some food | pot, oven, sauce pan | Llama-3-8b-Instruct | pot (54%), others(0%) |
| Gemma-2-9b-it | pot (52%), others(0%) | ||
| Wash dishes with dishwasher | dishwasher, dish soap, bowl, plate, drinking glass, | Llama-3-8b-Instruct | dishwasher (100%), plate (38%), others (0%) |
| fork, cup, spoon, pot, knife | Gemma-2-9b-it | dishwasher (100%), dish soap(18%), plate (39%), others (0%) | |
| Wash clothes | washing machine, soap, pant, shirt, dress, jacket | Llama-3-8b-Instruct | Washing machine (100%), soap(64%), others (0%) |
| Gemma-2-9b-it | washing machine (100%), soap(62%), others (0%) |
Appendix H Hallucination Probing Prompts Design and Distribution
H.1 Base Prompt Format and Structure